# A bound on the quantum value of all compiled nonlocal games

### v1

@@ Abstract.
A compiler introduced by Kalai et al. (STOC’23) converts any nonlocal game into an interactive protocol with a single computationally-bounded prover. Although the compiler is known to be sound in the case of classical provers, as well as complete in the quantum case, quantum soundness has so far only been established for special classes of games.
@@
In this work, we establish a quantum soundness result for all compiled two-player nonlocal games. In particular, we prove that the quantum commuting operator value of the underlying nonlocal game is an upper bound on the quantum value of the compiled game. Our result employs techniques from operator algebras in a computational and cryptographic setting to establish information-theoretic objects in the asymptotic limit of the security parameter. It further relies on a sequential characterization of quantum commuting operator correlations which may be of independent interest.
1. Introduction
@@
A nonlocal game consists of two (or more) non-communicating players interacting with a referee. In the game, the referee samples a question for each player, to which each player replies with an answer. The referee decides if the players win or lose based on the tuple of questions and answers. Since communication is not permitted between players, each player has no information about the questions given to the other players, nor do they know the answers provided to the referee by the other players. Nevertheless, the description of the game is known to the players ahead of time, allowing them to strategize and maximize their probability of winning the game. The classical value $\omega_{c}(\mathcal{G})$ of a nonlocal game $\mathcal{G}$ is the maximum winning probability of players with classical resources (such as shared randomness). On the other hand, the quantum value $\omega_{q}(\mathcal{G})$ represents the maximum winning probability of quantum players sharing a finite amount of quantum resources (such as entangled quantum states, like EPR pairs). In the quantum setting, the no-communication assumption can either be modeled by (i) spatially separating the player so that they act on tensor product subsystems or (ii) requiring that the players’ actions commute on the joint system. While these two conditions are equivalent when the state space is finite dimensional, they are not when one allows infinite-dimensional systems. This motivates the study of the (quantum) commuting operator value $\omega_{qc}(\mathcal{G})$, which denotes the maximum winning probability over strategies where the measurement operators of the players commute.
@@
Nonlocal games can be viewed as variants of Bell scenarios [Bel64] and been successful in advancing our understanding of entanglement. In particular, have been a productive framework for separating various sets of correlations arising from various physical models. Notably, they have been used to discern the sets of classical and quantum correlations $C_{c}\subsetneq C_{q}$ [CHSH69], the sets of quantum correlations and their closure (the quantum approximable correlations) $C_{q}\subsetneq C_{qa}$ [Slo17], and the sets of quantum approximable and commuting operator correlations $C_{qa}\subsetneq C_{qc}$ [JNV+21]. Along the way, nonlocal games became an important topic in complexity theory, through their connection to multiprover interactive proofs [CHTW04]. More recently, nonlocal games have enabled new protocols to certify quantum computation in the two-prover setting [RUV13, CGJV19, Gri19, JNV+21]. A fundamental question in this area is whether two non-communicating provers are really necessary to build such protocols. The single-prover setting, where a verifier interacts with a single computationally-bounded prover, is both theoretically appealing and practically motivated since the non-communication assumption can be difficult to enforce.
@@
Motivated by this issue, Kalai, Lombardi, Vaikuntanathan, and Yang (KLVY) proposed a generic procedure to transform any nonlocal game into a single-prover protocol, replacing the no-communication assumption between players with a computational assumption on the prover [KLVY23]. For instance, the KLVY compiler translates a two-player game into a four-round game with a single player (prover) and a referee (verifier). Questions are asked and answered sequentially, rather than in parallel, and the leaking of information to the next round is prevented by cryptographic assumptions as signaling in the other direction is prevented by the temporal order of the game. To achieve the desired functionality the construction employs a quantum homomorphic encryption scheme with classical messages [Mah18, Bra18]. This results in a quantum polynomial time (QPT) assumption on the prover, as any prover with greater computational resources could break the security of the QHE scheme, and leak information about the encrypted questions.
@@
In addition to outlining the compilation procedure, Kalai et al. established classical soundness and quantum completeness of the compiler, meaning that while classical provers cannot exceed the classical value of the corresponding nonlocal game, quantum provers can achieve the quantum value, in the asymptotic limit where the security parameter $\lambda$ of the underlying quantum homomorphic encryption scheme tends to infinity. In particular this implies that any nonlocal game with $\omega_{c}<\omega_{q}$ can be converted into a protocol that classically verifies quantum advantage.
@@
In this work we consider the problem of establishing soundness of the compiler in the quantum setting. Quantum soundness means that the quantum value of the compiled game is suitably bounded above by a quantum value for the original game. A series of subsequent works established upper bounds on the quantum compiled value for the CHSH game [NZ23], the class of XOR games [CMM+24, BVB+24], tilted-CHSH scenarios [MPW24], and for self-tests on Pauli measurements on maximally entangled states [MNZ24]. This enables an alternative and conceptual attractie way to perform verification of BQP and QMA computations with a classical verifier. Despite this progress, a general bound on the quantum value that applies to all compiled nonlocal games remained elusive.
1.1. Main result and techniques
@@
In this work, we show that the quantum value of any compiled two-player nonlocal game is bounded by the quantum commuting operator value of the underlying nonlocal game.
Theorem (Informal).
@@
For large enough security parameter, no quantum polynomial time prover can win the compiled game with probability exceeding the quantum commuting operator value of the game by any constant.
@@
In other words, we show that for any nonlocal game $\mathcal{G}$,
$\omega_{q,\max}(\mathcal{G}_{\textnormal{comp}})\leq\omega_{qc}(\mathcal{G}),$
@@
where $\omega_{q,\max}(\mathcal{G}_{\textnormal{comp}})$ denotes the supremum of the winning probability of any quantum polynomial time prover against the compiled game $\mathcal{G}_{\textnormal{comp}}$, in the limit that the security parameter tends to infinity. The definition of $\mathcal{G}_{\textnormal{comp}}$ is given in Definition 4.2 whereas $\omega_{q,\max}(\mathcal{G}_{\textnormal{comp}})$ is defined in Definition 4.4. The formal results are given in Theorems 6.1 and 6.2. Our theorem establishes quantum commuting operator soundness for all compiled two-player nonlocal games.
Our main result relies on two key ideas. First, we present a new characterization of quantum commuting operator correlations in terms of strategies for a sequential game, which serves as an idealization of the compiled game. We show that quantum sequential strategies that satisfy a certain strong non-signaling property give rise to quantum commuting operator correlations. We believe this alternative characterization of commuting operator correlations may be of independent interest.
Our second idea concerns the observation that the strong non-signaling property required in the idealized sequential game to obtain commuting operator correlations is obtained only in the limit of the security parameter tending to infinity. To make sense of compiled quantum strategies with respect to this limit we incorporate mathematical tools from operator algebras. Specifically we show that the essential part of the prover’s strategies can be captured by states on a universal $C^{*}$-algebra and use a compactness argument to prove the existence of a limiting state. The limiting state can be then shown to precisely satisfy the aforementioned non-signaling like property, even though for every finite level of the security parameter it only holds for efficient observables. This approach is reminiscent of proofs of completeness for noncommutative optimization hierarchies such as the NPA hierarchy [NPA08].
1.2. Comparison with prior work

### v3

@@ Abstract
A cryptographic compiler introduced by Kalai, Lombardi, Vaikuntanathan, and Yang (STOC’23) converts any nonlocal game into an interactive protocol with a single computationally bounded prover. Although the compiler is known to be sound in the case of classical provers and complete in the quantum case, quantum soundness has so far only been established for special classes of games.
@@
In this work, we establish a quantum soundness result for all compiled two-player nonlocal games. In particular, we prove that the quantum commuting operator value of the underlying nonlocal game is an upper bound on the quantum value of the compiled game. Our result employs techniques from operator algebras in a computational and cryptographic setting to establish information-theoretic objects in the asymptotic limit of the security parameter. It further relies on a sequential characterization of quantum commuting operator correlations, which may be of independent interest.
1 Introduction
@@
A nonlocal game consists of two (or more) cooperative players that interact with a referee. In the game, the referee samples a question for each player, to which each player replies with an answer. The referee decides if the players win or lose based on the tuple of questions and answers. Communication is not permitted between players, hence each player has no information about the questions given to the other players, nor do they know the answers provided to the referee by the other players. Nevertheless, the description of the game is known to the players ahead of time, allowing them to strategize and maximize their probability of winning the game. The classical value $\omega_{c}(\mathcal{G})$ of a nonlocal game $\mathcal{G}$ is the maximum winning probability of classical players, while the quantum value $\omega_{q}(\mathcal{G})$ denotes the maximum winning probability of quantum players sharing a finite amount of quantum resources (such as entangled quantum states, like EPR pairs). In the quantum setting, the no-communication assumption can either be modeled by (i) spatially separating the players so that they act on tensor product subsystems or (ii) requiring that the players’ actions commute on the joint system. While these two conditions are equivalent when the state space is finite-dimensional, they are inequivalent in infinite dimensions. The quantum commuting operator value $\omega_{qc}(\mathcal{G})$ denotes the maximum winning probability over strategies where the players’ measurement operators commute.
@@
Nonlocal games have proved highly influential in quantum information. They have significantly advanced and operationalized our understanding of entanglement [4], and offered a productive framework for separating correlations arising from various physical models. Notably, they have been used to discern the classical and quantum correlations $C_{c}\subsetneq C_{q}$ [9], the quantum correlations and their closure $C_{q}\subsetneq C_{qa}$ [42], and the latter from the commuting operator correlations $C_{qa}\subsetneq C_{qc}$ as a consequence of $\mathsf{MIP}^{*}=\mathsf{RE}$ [22]. Along the way, they became an important topic in complexity theory, through their connection to multiprover interactive proofs [10], and they allow certifying quantum computation in the two-prover setting [40, 12, 20, 22].
1.1 Background on compiled nonlocal games
@@
A fundamental question in this area is whether two non-communicating provers are really necessary to build such protocols. The single-prover setting, where a verifier interacts with a single computationally-bounded prover, is both theoretically appealing and practically motivated since the no-communication assumption can be difficult to enforce. A breakthrough by Mahadev showed that quantum computations can be verified in this setting [25]. To establish a conceptual connection between the two worlds, Kalai, Lombardi, Vaikuntanathan, and Yang (KLVY) proposed a generic procedure to transform any nonlocal game Fig. 1(a) into a single-prover protocol, called the compiled nonlocal game Fig. 1(b), replacing the no-communication assumption between players with a computational assumption on the prover [24]. For instance, the KLVY compiler translates a two-player game into a four-round game with a single player (prover) and a referee (verifier). Questions are asked and answered sequentially, rather than in parallel, and the leaking of information to the next round is prevented by cryptographic assumptions. To achieve the desired functionality the construction employs a quantum homomorphic encryption (QHE) scheme with classical messages [26, 6], which is used to encrypt the first question (and consequently the answer) of the prover, whereas the second question is sent in plain. This results in a quantum polynomial time (QPT) assumption on the prover, as any prover with greater computational power could break the security of the QHE scheme, and gain information about the encrypted questions.
@@
In addition to outlining the compilation procedure, [24] established classical soundness and quantum completeness of the compiler. This means that while classical provers cannot exceed the classical value of the corresponding nonlocal game, quantum provers can even achieve the quantum value, in the asymptotic limit where the security parameter $\lambda$ (of the underlying quantum homomorphic encryption scheme) tends to infinity. This implies that any nonlocal game with $\omega_{c}<\omega_{q}$ can be converted into a protocol that classically verifies quantum advantage.
(a) Two-player nonlocal game
(b) Compiled nonlocal game [24]
(c) Sequential game without cryptography
Figure 1: Fig. 1(a) depicts a conventional nonlocal game between a verifier $\mathsf{V}$ and non-communicating computationally unbounded quantum provers $\mathsf{A}$ and $\mathsf{B}$. Fig. 1(b) depicts a compiled game with verifier $\mathsf{V}$ and a single computationally bounded quantum prover $\mathsf{P}$. The security of the encryption scheme ensures that post measurement states following the first round of communication are computationally indistinguishable. Fig. 1(c) depicts a sequential game, which is an idealized compiled game, with a verifier $\mathsf{V}$ and a single quantum prover $\mathsf{P}$. In a sequential game we assume that all post-measurement states after the first round of communication are identical.
@@
However, establishing quantum soundness, i.e., an upper bound on the success probability that can be achieved by QPT strategies for the compiled game, has proven to be a more difficult task. Recent works have achieved a bound only for special classes of games, such as the CHSH game [33], the class of XOR games [15, 2], tilted-CHSH scenarios [28], and for self-tests on Pauli measurements on maximally entangled states [30]. As a consequence of these works, compiled nonlocal games paved the way for a conceptually elegant and modular way to perform verification of BQP and QMA computations with a classical verifier. Despite this progress, establishing a general bound on the quantum value that applies to all compiled nonlocal games remained elusive.
1.2 Main results
@@
In this work, we make progress on this fundamental open problem by showing that the quantum value of the compiled game is always upper-bounded by the quantum commuting operator value of the nonlocal game, for any arbitrary two-player nonlocal game. To achieve this we also introduce some new conceptual tools.
Theorem 1.1.
@@
For large enough security parameter $\lambda$, no QPT strategy can win the compiled game with probability exceeding the quantum commuting operator value of the game by any constant.
@@
In other words, we show that for any two-player nonlocal game $\mathcal{G}$ and for any quantum polynomial-time (QPT) strategy $S$ for the compiled game $\mathcal{G}_{\textnormal{comp}}$,
$\limsup_{\lambda\to\infty}\omega_{\lambda}(\mathcal{G}_{\textnormal{comp}},S)\leq\omega_{qc}(\mathcal{G}),$
@@
where $\omega_{\lambda}(\mathcal{G}_{\textnormal{comp}},S)$ denotes the winning probability of the strategy $S$ as a function of the security parameter $\lambda$ (Definition 4.2). Thus, our theorem establishes quantum commuting operator soundness for all compiled two-player nonlocal games. The proof of Theorem 1.1 follows directly from Theorems 6.1 and 6.2.
At a technical level, our proof combines methods from operator algebras in the context of cryptographic protocols. Our proof builds on two technical ingredients, that might be of independent interest:
  1. (i)
    a sequential characterization of quantum commuting operator correlations, motivated in an idealization of the security guarantee offered by the KLVY compiler Fig. 1(c), and
  2. (ii)
    the idea of analyzing cryptographic protocols with computational security by taking the limit of the security parameter to infinity to obtain a protocol that offers information-theoretic security.
Our characterization generalizes a result in [34] that describes the spatial quantum correlations in terms of so-called “quansal” correlations. It also solves an open problem in the context of relating Bell and prepare-and-measure scenarios and strategies, characterizing the image of $C_{qc}$ under the mapping defined in [46]. The characterization of commuting operator correlations was also established independently in the context of steering in [1].